# When in Doubt, Consult: Expert Debate for Sexism Detection via Confidence-Based Routing

### v2

@@ Abstract
Sexist content online increasingly appears in subtle, context-dependent forms that evade traditional detection methods. Its interpretation often depends on overlapping linguistic, psychological, legal, and cultural dimensions, which produce mixed and sometimes contradictory signals, even in annotated datasets. These inconsistencies, combined with label scarcity and class imbalance, result in unstable decision boundaries and cause fine-tuned models to overlook subtler, underrepresented forms of harm. Together, these limitations point to the need for a design that explicitly addresses the combined effects of (i) underrepresentation, (ii) noise, and (iii) conceptual ambiguity in both data and model predictions. To address these challenges, we propose a two-stage framework that unifies (i) targeted training procedures to adapt supervision to scarce and noisy data with (ii) selective, reasoning-based inference to handle ambiguous or borderline cases. Our training setup applies class-balanced focal loss, class-aware batching, and post-hoc threshold calibration to mitigate label imbalance and noisy supervision. At inference time, a dynamic routing mechanism classifies high-confidence cases directly and escalates uncertain instances to a novel Collaborative Expert Judgment (CEJ) module, which prompts multiple personas and consolidates their reasoning through a judge model. Our approach achieves state-of-the-art results across several benchmarks, with F1 gains of +4.48% and +1.30% on EDOS Tasks A and B, respectively, and a +2.79% improvement in ICM on EXIST 2025 Task 1.1.
@@ 1 Introduction
Sexist content on social media is increasingly pervasive, often appearing in subtle, sarcastic, or context-dependent forms that evade traditional classification methods. From seemingly innocuous stereotypes to overt hate speech, such content contributes to the emotional and psychological harm that disproportionally affects women abercrombie-etal-2023-resources.
@@
These diverse forms of sexism are shaped by distinct yet overlapping factors: linguistic, psychological, legal, and cultural, among others. As a result, both human annotators and automated systems must contend with mixed and sometimes contradictory signals: a message may appear harmless to a general audience, yet be perceived as harmful by a legal expert, a psychologist, or someone with lived experience of gender-based discrimination davani2022disagreement.
@@
This multidimensionality creates several challenges for annotation and automatic detection. Even the largest and most reliable datasets edos23; exist25 are typically small, highly imbalanced, and marked by substantial annotator disagreement not only due to ambiguity in individual examples, but also because different annotators may implicitly prioritize different criteria (e.g., linguistic form, psychological impact, cultural normativity, or legal definitions) khan2025explaining. These mixed signals can overshadow the representation of less explicit but equally damaging forms of sexism, introducing noise and instability into both the data and the models. At the same time, large language models (LLMs) are pretrained on vast, heterogeneous data sources, absorbing social norms, stereotypes, which are inconsistently represented, and at times in contradiction with human feedback. As a result, fine-tuning these large models on small, noisy sexism datasets is often insufficient to resolve these conflicts, and may inadvertently reinforce dominant or explicit patterns while neglecting subtler cases khan2025explaining. These limitations underscore the need for models that explicitly account for the combined effect of underrepresentation, noise, and conceptual ambiguity in both data and predictions.
@@
To address these challenges, we introduce a two-stage framework that combines (i) targeted training procedures to adapt supervision to scarce and noisy data, with (ii) selective, reasoning-based inference to handle ambiguous cases. This joint design enables the specialized model to make efficient decisions when possible, while engaging with finer contextual reasoning when necessary. During (i) training, we mitigate long-tailed label distributions and inconsistent annotation through three well-known training strategies. These have never been adapted into an integrated framework for such tasks: the class-balanced focal loss (CB-Focal) (ce), which amplifies the contribution of minority and difficult examples, and the class-aware batching, which ensures uniform exposure to underrepresented classes during optimization. We further employ post-hoc threshold calibration to stabilize the decision boundaries of the specialist classifier and improve robustness under noisy supervision. At (ii) inference time, we deploy confidence-aware dynamic routing, which filters the model predictions by difficulty: high-confidence samples are directly classified, while uncertain cases are escalated to a Collaborative Expert Judgment (CEJ) module. The CEJ module implements a structured reasoning pipeline composed of multiple personas, such as a linguist, a gender expert, a psychologist, and an everyday speaker, each prompted to analyze the same input from distinct perspectives. Their deliberations are subsequently synthesized by a judge model, which integrates these perspectives into a final, interpretable decision (Figure 1). The aim is not to faithfully replicate human interaction, but rather to approximate its structure to activate the reasoning and knowledge that instruction-tuned LLMs already encode, yet rarely employ in single-step classification, as evidenced by the limited predictive power of zero-shot baselines edos23; exist25; khan2025explaining.
@@ 2 Related Work
Annotator disagreement, often shaped by cultural perspectives, complicates sexism detection davani2022disagreement. Initiatives like the EXIST 2024 shared task encouraged methods robust to inconsistent annotations. Successful approaches leveraged transformer-based models such as DistilBERT sanh2019distilbert, DeBERTa he2020deberta, RoBERTa liu2019roberta, and hybrid ensembles that utilized data augmentation and multi-task learning. When it comes to LLMs, previous studies have employed adversarial evaluations and data augmentation to enhance detection accuracy. For example, samory2021call highlighted standard model vulnerabilities through psychologically grounded adversarial datasets, emphasizing the need for psychologically informed robustness. Similarly, khan2025explaining addressed annotator disagreements through definition-driven data augmentation and ensemble methods, improving data-level robustness.
Beyond explicit abuse, a growing challenge lies in detecting nuanced or implicit forms of hate. The work by zeng2025sheeps shows that LLMs often fail to recognize metaphorical expressions of hate speech, highlighting substantial gaps in current moderation capabilities. Their findings emphasize that even state-of-the-art models struggle when hateful intent is obscured through indirect language, a phenomenon that is particularly relevant for sexist discourse. Building on this, the authors in fasching2025model demonstrate that hate speech detection outcomes vary significantly across LLM-based systems, underscoring how moderation decisions remain model-dependent and inconsistent. Our work directly responds to these challenges by integrating confidence-aware routing and multi-expert arbitration, ensuring that inconsistent or ambiguous cases are addressed through structured reasoning and synthesis.
@@
More recent work integrates multi-agent architectures, expert simulations, and human-in-the-loop reasoning to improve transparency and social norm alignment. While expert persona-driven prompting strategies, such as those in xu2023expertprompting; long2024multi, have advanced interpretability, our method surpasses static personas by introducing dynamic, structured, and confidence-aware interactions among multiple expert roles, capturing richer contextual nuances. Moreover, our framework integrates a dedicated judge model to arbitrate conflicts, providing nuanced decisions beyond simple voting mechanisms.
Emerging research demonstrated the effectiveness of multi-expert deliberation in LLM-based decision-making. lu-etal-2024-triageagent introduce a multi-agent framework for clinical triage where LLM agents assume distinct roles, engage in multi-round discussions with self-confidence scoring, and iteratively revise their assessments until reaching consensus. Similarly, long2024multi simulate multiple domain experts within a single LLM, aggregate their responses, and select the best one. Both approaches show that structured discussion among diverse perspectives yields more reliable outputs than single-viewpoint reasoning, as cross-perspective validation mitigates individual biases and enables principled conflict resolution before final judgment. Our framework extends this paradigm to sexism detection, where the inherently subjective nature of the task makes perspectival diversity essential.
Another notable study that effectively leverages expert agents is Expert-Token-Routing chai2024expert, which adopts a router-based approach similar to that of jiang2023llm. In this method, a router model is trained to direct each question or instruction to a specialized expert, where the expert LLM is responsible for generating the next token. Their approach encodes expert LLMs as special tokens within the vocabulary of the meta-LLM, drawing inspiration from ToolkenGPT hao2023toolkengpt.
Our progressive prompt refinement aligns with iterative human-in-the-loop methods discussed in shah2025prompt, systematically enhancing logical coherence and prompting effectiveness. In addition, we preserve the logical structure and enhance the prompt formatting for effective model reasoning as it is shown by li2025llms’s experiments.
@@ 3 Methodology
We introduce a two-stage framework for sexism detection. The framework combines a (i) specialized neural model with (ii) a selective reasoning component. The specialized model is fine-tuned using adapted methodologies to improve robustness under both data imbalance and noisy supervision. At inference, a confidence-aware routing mechanism selects high-confidence instances for the specialised model or escalates uncertain ones to a Collaborative Expert Judgment (CEJ) module, where multiple personas deliberate and a judge model produces the final decision (Figure 2).
@@ 3.1 Domain-Tuned Model
We fine-tune a small instruction-tuned LLM (LLaMA-3.2-3B) as a domain-specialized classifier. To handle severe class imbalance, we integrate three complementary strategies: Class-Balanced Cross-Entropy (CB-CE) (ce) to reweight loss by effective sample count, Focal Loss (ce) to emphasize difficult minority-class instances, and Class-Aware Batch Sampling (CAB) (henning23) to ensure balanced gradient contributions. For binary classification, we apply temperature scaling (guo2017calibration) and tune a decision threshold on the development set; for multi-class classification, we use CB-Focal loss given the more severe imbalance.
@@ 3.2 Confidence-Aware Routing
At this stage, all the low-confidence predictions from the specialised models are escalated to the collaborative reasoning method for a multi-perspective analysis. Specifically, we adopt a selective classification rule (elyaniv2010foundations) given an input $x$, the model produces a predicted label $\hat{y}_{s}$ with the associated confidence score $c_{s}(x)$:
@@
In multi-class settings, low confidence may not fully capture uncertainty, particularly when the model is uncertain between its top two predictions. We define the margin between the two most probable labels as $m(x)=p_{(1)}(x)-p_{(2)}(x)$ where $p_{(1)}(x)$ and $p_{(2)}(x)$ denote the largest and second-largest posterior probabilities. An instance is then routed to CEJ if:
@@ 3.3 Collaborative Expert Judgment (CEJ)
Prior research on abusive and sexist language highlights the importance of combining expert knowledge with lived experience sap2020socialbiasframes; davani2022disagreement. Inspired by this, we define a set of personas representing diverse interpretive perspectives. Expert roles, such as linguists, psychologists, gender studies scholars, and legal professionals have been shown to provide critical perspectives on language use, structural bias, emotional impact, and legal norms lazar2005feminist; vidgen2019challenges. To still capture non-expert and experiential perspectives, we also incorporate a sexism victim persona, reflecting direct experience of harm, and a layperson persona, capturing everyday interpretations that frequently underlie annotator disagreement (Table A3).
It is important to emphasize that the aim of this multi-persona prompting is not to simulate faithful and exhaustive expert panels, but rather to elicit and organize the contextual, normative, and experiential knowledge already encoded within large language models. By inducing the model to articulate these diverse perspectives, the CEJ module encourages the application of domain understanding that often remains latent in single-step classification.
# Variant Model ICM ICM-Norm F1_YES F1 Zero-shot Baselines 1 GPT-5.2 0.3408 0.6705 0.7043 0.7349 2 Qwen2.5-72B-Instruct 0.3765 0.6892 0.7112 0.7384 3 LLaMA-3.3-70B-Instruct 0.3002 0.6509 0.6868 0.7145 4 Cogito-70B 0.3656 0.6837 0.7109 0.7348 5 Cogito-70B (reasoning) 0.3772 0.6896 0.7143 0.7384 EXIST 2025 Task 1.1 Leaderboard 6 BERT nowakowski2025exist 0.5727 0.7878 0.7802 7 XLM-RoBERTa pan2025exist 0.5799 0.7915 0.7824 8 Dual-Transformer Fusion Network khan2025exist 0.5806 0.7918 0.7837 9 DeepSeek-R1-Distill-Llama-8B villarreal2025exist 0.6127 0.8079 0.7945 10 Ensemble Approach c:22 0.6249 0.8141 0.7991 11 XLM-RoBERTa c:21 0.6297 0.8165 0.7996 12 LLaMA-3.1-8B-Instruct tian2025marioexist2025simple 0.6774 0.8405 0.8167 Ours 13 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.5709 0.7913 0.7596 0.7934 14 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.6595 0.8315 0.8089 0.8257 15 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.6831 0.8433 0.8191 0.8348 16 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.6963 0.8500 0.8233 0.8389
Table 1: Results on the EXIST 2025 Task 1.1 test set with zero-shot baselines, leaderboard submissions, and our approach. "Q" and "L" denote the models used for personas debate (Qwen2.5-72B-Instruct and LLaMA-3.3-70B, respectively), and "C" refers to the judge model Cogito-70B (reasoning mode). Bold = best score, underline = runner-up.
@@ 3.3.1 Structured Debate and Judge
We operationalize the CEJ module as a multi-stage panel, in which each persona offers a distinct interpretive perspective on the input. The process comprises four sequential stages, designed to elicit, interrogate, and synthesize domain-relevant knowledge:
Initial Opinions. Each persona independently analyzes the input, providing an initial classification, justification, with a confidence score. This initial stage establishes diverse baseline perspectives grounded in each persona’s unique expertise.
Structured Debate. Personas are subsequently exposed to all initial opinions and engage in critical evaluation of their peers’ reasoning. During this deliberation, each persona must: (i) engage with at least one other perspective, explicitly agreeing or disagreeing and offering supporting rationale; (ii) revise their own stance if confronted with compelling counter-arguments; and (iii) reassess the input through the interpretive lens of other personas, re-evaluating both intent and target. This process results in a revised classification, an updated justification, and an adjusted confidence score for each persona.
Summarization. A dedicated summarization agent condenses the full debate into a concise synthesis highlighting the main arguments, points of consensus, and unresolved disagreements.
Final Judgment. Finally, the judge model synthesizes all available evidence, i.e., the original input, both initial and revised persona opinions, and the debate summary, to produce a final adjudication. Prompted to act impartially, the judge delivers a definitive label, comprehensive justification, and confidence score that resolves any lingering disagreements.
# Variant Model Task A Task B Task C Zero-shot Baselines 1 GPT-5.2 0.7306 0.4543 0.3381 2 Qwen-2.5-72B-Instruct 0.6190 0.4246 0.2576 3 LLaMA-3.3-70B-Instruct 0.6421 0.4297 0.2775 4 Cogito-70B 0.6106 0.4214 0.2683 5 Cogito-70B (reasoning) 0.6332 0.4325 0.2826 SemEval 2023 Task 10.A Leaderboard 6 DeBERTa-v3-large + twHIN-BERT-large zhou-2023-pinganlifeinsurance 0.8746 7 RoBERTa-Large + ELECTRA 0.8740 0.7203 0.5487 8 DeBERTa Ensemble 0.8740 Data Augmentation and Ensemble Methods 9 SEFM zhong-etal-2023-uirisc 0.8538 0.6619 0.4641 10 QCon feely-etal-2023-qcon 0.8400 0.6400 0.4700 11 HULAT segura-bedmar-2023-hulat-semeval 0.8298 0.5877 0.4458 12 CSE khan2025explaining 0.8819 0.7243 0.5639 13 DDA khan2025explaining 0.8769 0.7277 0.6018 14 PaLM Ensemble 0.7326 Ours 15 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.7824 0.5941 0.3717 16 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.8986 0.6279 0.4189 17 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.9195 0.7324 0.5842 18 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.9214 0.7421 0.5904

### v3

@@ Abstract
Online sexism increasingly appears in subtle, context-dependent forms that evade traditional detection methods. Its interpretation often depends on overlapping linguistic, psychological, legal, and cultural dimensions, which produce mixed and sometimes contradictory signals in annotated datasets. These inconsistencies, combined with label scarcity and class imbalance, result in unstable decision boundaries and cause fine-tuned models to overlook subtler, underrepresented forms of harm. To address these challenges, we propose a two-stage framework that unifies (i) targeted training procedures to better regularize supervision to scarce and noisy data with (ii) selective, reasoning-based inference to handle ambiguous or borderline cases. First, we stabilize the training combining class-balanced focal loss, class-aware batching, and post-hoc threshold calibration, strategies for the firs time adapted for this domain to mitigate label imbalance and noisy supervision. Second, we bridge the gap between efficiency and reasoning with a a dynamic routing mechanism that distinguishes between unambiguous instances and complex cases requiring a deliberative process. This reasoning process results in the novel Collaborative Expert Judgment (CEJ) module which prompts multiple personas and consolidates their reasoning through a judge model. Our approach outperforms existing approaches across several public benchmarks, with F1 gains of +4.48% and +1.30% on EDOS Tasks A and B, respectively, and a +2.79% improvement in ICM on EXIST 2025 Task 1.1.
@@ 1 Introduction
Sexist content on social media is increasingly pervasive, often appearing in subtle, sarcastic, or context-dependent forms that evade traditional classification methods. From seemingly innocuous stereotypes to overt hate speech, such content contributes to the emotional and psychological harm that disproportionally affects women Abercrombie et al. (2023).
@@
These diverse forms of sexism are shaped by distinct yet overlapping factors: linguistic, psychological, legal, and cultural, among others. As a result, both human annotators and automated systems must contend with mixed and sometimes contradictory signals: a message may appear harmless to a general audience, yet be perceived as harmful by a legal expert, a psychologist, or someone with lived experience of gender-based discrimination Mostafazadeh Davani et al. (2022).
@@
This multidimensionality creates several challenges for annotation and automatic detection. Even the largest and most reliable datasets Kirk et al. (2023); Plaza et al. (2025) are typically small, highly imbalanced, and marked by substantial annotator disagreement not only due to ambiguity in individual examples, but also because different annotators may implicitly prioritize different criteria (e.g., linguistic form, psychological impact, cultural normativity, or legal definitions) Khan et al. (2025a). These mixed signals can overshadow the representation of less explicit but equally damaging forms of sexism, introducing noise and instability into both the data and the models. At the same time, large language models (LLMs) are pretrained on vast, heterogeneous data sources, absorbing social norms, stereotypes, which are inconsistently represented, and at times in contradiction with human feedback. As a result, fine-tuning these large models on small, noisy sexism datasets is often insufficient to resolve these conflicts, and may inadvertently reinforce dominant or explicit patterns while neglecting subtler cases Khan et al. (2025a). These limitations underscore the need for models that explicitly account for the combined effect of underrepresentation, noise, and conceptual ambiguity in both data and predictions.
@@
To address these challenges, we introduce a two-stage framework that combines (i) targeted training procedures to adapt supervision to scarce and noisy data, with (ii) selective, reasoning-based inference to handle ambiguous cases. This joint design enables the specialized model to make efficient decisions when possible, while engaging with finer contextual reasoning when necessary. During (i) training, we mitigate long-tailed label distributions and inconsistent annotation through three well-known training strategies. These have never been adapted into an integrated framework for such tasks: the class-balanced focal loss (CB-Focal) (Cui et al., 2019), which amplifies the contribution of minority and difficult examples, and the class-aware batching, which ensures uniform exposure to underrepresented classes during optimization. We further employ post-hoc threshold calibration to stabilize the decision boundaries of the specialist classifier and improve robustness under noisy supervision. At (ii) inference time, we deploy confidence-aware dynamic routing, which filters the model predictions by difficulty: high-confidence samples are directly classified, while uncertain cases are escalated to a Collaborative Expert Judgment (CEJ) module. The CEJ module implements a structured reasoning pipeline composed of multiple personas, such as a linguist, a gender expert, a psychologist, and an everyday speaker, each prompted to analyze the same input from distinct perspectives. Their deliberations are subsequently synthesized by a judge model, which integrates these perspectives into a final, interpretable decision (Figure 1). The aim is not to faithfully replicate human interaction, but rather to approximate its structure to activate the reasoning and knowledge that instruction-tuned LLMs already encode, yet rarely employ in single-step classification, as evidenced by the limited predictive power of zero-shot baselines Kirk et al. (2023); Plaza et al. (2025); Khan et al. (2025a).
@@ 2 Related Work
Annotator disagreement, often shaped by cultural perspectives, complicates sexism detection Mostafazadeh Davani et al. (2022). Initiatives like the EXIST 2024 shared task encouraged methods robust to inconsistent annotations. Successful approaches leveraged transformer-based models such as DistilBERT Sanh et al. (2019), DeBERTa He et al. (2020), RoBERTa Liu et al. (2019), and hybrid ensembles that utilized data augmentation and multi-task learning. When it comes to LLMs, previous studies have employed adversarial evaluations and data augmentation to enhance detection accuracy. For example, Samory et al. (2021) highlighted standard model vulnerabilities through psychologically grounded adversarial datasets, emphasizing the need for psychologically informed robustness. Similarly, Khan et al. (2025a) addressed annotator disagreements through definition-driven data augmentation and ensemble methods, improving data-level robustness.
Beyond explicit abuse, a growing challenge lies in detecting nuanced or implicit forms of hate. The work by Zeng et al. (2025) shows that LLMs often fail to recognize metaphorical expressions of hate speech, highlighting substantial gaps in current moderation capabilities. Their findings emphasize that even state-of-the-art models struggle when hateful intent is obscured through indirect language, a phenomenon that is particularly relevant for sexist discourse. Building on this, the authors in Fasching and Lelkes (2025) demonstrate that hate speech detection outcomes vary significantly across LLM-based systems, underscoring how moderation decisions remain model-dependent and inconsistent. Our work directly responds to these challenges by integrating confidence-aware routing and multi-expert arbitration, ensuring that inconsistent or ambiguous cases are addressed through structured reasoning and synthesis.
@@
More recent work integrates multi-agent architectures, expert simulations, and human-in-the-loop reasoning to improve transparency and social norm alignment. While expert persona-driven prompting strategies, such as those in Xu et al. (2023); Long et al. (2024), have advanced interpretability, our method surpasses static personas by introducing dynamic, structured, and confidence-aware interactions among multiple expert roles, capturing richer contextual nuances. Moreover, our framework integrates a dedicated judge model to arbitrate conflicts, providing nuanced decisions beyond simple voting mechanisms.
Emerging research demonstrated the effectiveness of multi-expert deliberation in LLM-based decision-making. Lu et al. (2024) introduce a multi-agent framework for clinical triage where LLM agents assume distinct roles, engage in multi-round discussions with self-confidence scoring, and iteratively revise their assessments until reaching consensus. Similarly, Long et al. (2024) simulate multiple domain experts within a single LLM, aggregate their responses, and select the best one. Both approaches show that structured discussion among diverse perspectives yields more reliable outputs than single-viewpoint reasoning, as cross-perspective validation mitigates individual biases and enables principled conflict resolution before final judgment. Our framework extends this paradigm to sexism detection, where the inherently subjective nature of the task makes perspectival diversity essential.
Another notable study that effectively leverages expert agents is Expert-Token-Routing Chai et al. (2024), which adopts a router-based approach similar to that of Jiang et al. (2023). In this method, a router model is trained to direct each question or instruction to a specialized expert, where the expert LLM is responsible for generating the next token. Their approach encodes expert LLMs as special tokens within the vocabulary of the meta-LLM, drawing inspiration from ToolkenGPT Hao et al. (2023).
Our progressive prompt refinement aligns with iterative human-in-the-loop methods discussed in Shah (2025), systematically enhancing logical coherence and prompting effectiveness. In addition, we preserve the logical structure and enhance the prompt formatting for effective model reasoning as it is shown by Li et al. (2025)’s experiments.
@@ 3 Methodology
We introduce a two-stage framework for sexism detection. The two stages are based on the core premise that not all instances require complex reasoning. And while many can be better resolved by the specialist model, forcing indiscriminately an immediate decision on ambiguous cases leads to errors. We therefore implement a selective classification mechanism that acts as a gatekeeper between the (i) specialised classifier and the (ii) reasoning module. The specialized model is fine-tuned moving beyond standard cross-entropy minimization by integrating three regularization strategies designed to stabilize learning on the unstable decision boundaries caused by the severe class imbalance and annotator disagreement. Then, we couple this mechanism with a confidence-aware routing mechanism which selects low-confidence instances the needs to be escalated to the Collaborative Expert Judgment (CEJ) module, where multiple personas are designed to reason, deliberate, and produce the final decision (Figure 2).
@@ 3.1 Domain-Tuned Model
We first fine-tune a small instruction-tuned LLM as a domain-specialized classifier. To handle severe class imbalance, we integrate three complementary strategies: the Class-Balanced Cross-Entropy (CB-CE) (Cui et al., 2019), the Focal Loss (Cui et al., 2019), and the Class-Aware Batch Sampling (CAB) (Henning et al., 2023). With the CB-CE, rather than weighing samples by raw frequency, which can be misleading in long-tailed distributions, we calculate the effective number of samples to capture the diminishing marginal benefit of additional samples. This is combined with the Focal Loss objective, which dynamically down-weights easy negatives and focuses the gradient on hard examples, i.e., those that are misclassified or ambiguous. We further ensure consistent exposure to underrepresented categories through CAB, decoupling batch composition from the dataset’s natural distribution, so that gradient updates are not dominated by majority class signals. Finally, we apply post-hoc threshold calibration via temperature scaling (Guo et al., 2017), so that the model’s output probabilities accurately align with its epistemic uncertainty, preventing it from assigning high confidence to ambiguous cases where it lacks evidence. For binary classification, we apply temperature scaling and tune a decision threshold on the development set; for multi-class classification, we use CB-Focal loss given the more severe imbalance.
@@ 3.2 Confidence-Aware Routing
Following the initial classification, we employ a routing mechanism to handle instances where the specialised model exhibits high epistemic uncertainty, i.e., ambiguous examples. To systematically identify these uncertain predictions, we utilize a selective classification rule (El-Yaniv and others, 2010). For an input $x$, the model produces a predicted label $\hat{y}_{s}$ with the associated confidence score $c_{s}(x)$:
@@
In multi-class settings, reliance on raw confidence scores alone is insufficient because a model may often assign a relatively high probability to the top class while remaining deeply conflicted between two competing categories (e.g., Derogation vs. Animosity). We thus define the margin between the two most probable labels as $m(x)=p_{(1)}(x)-p_{(2)}(x)$ where $p_{(1)}(x)$ and $p_{(2)}(x)$ denote the largest and second-largest posterior probabilities. An instance is then routed to CEJ if:
@@ 3.3 Collaborative Expert Judgment (CEJ)
By acting as a gatekeeper, the aforementioned routing mechanism identify the instances where statistical pattern matching fails, and ensures that only the subset of difficult cases triggers the accurate but computationally expensive Collaborative Expert Judgment (CEJ) module. This reasoning module is based on prior research on abusive and sexist language which has widely highlighted the importance of combining expert knowledge with lived experience Sap et al. (2020); Mostafazadeh Davani et al. (2022). We therefore structure CEJ as a multi-perspective reasoning module based on a set of personas that act as distinct interpretive lenses. The ‘expert’ personas, such as linguists, psychologists, gender studies scholars, and legal professionals are chosen based on the literature showing that they are the main critical perspectives on language use, structural bias, emotional impact, and legal norms Lazar (2005); Vidgen and Derczynski (2019). Yet, to still capture non-expert and experiential perspectives, we incorporate a sexism victim persona, reflecting direct experience of harm, and a layperson persona, capturing everyday interpretations often responsible for annotator disagreement (Table A3).
It is worth noting that the aim of this multi-persona prompting is not to simulate faithful and exhaustive expert panels, but rather to elicit and organize the knowledge already encoded within LLMs. This combined approaches induces the CEJ module to trigger and controll the LLMs’ domain understanding that could otherwise remain latent in single-step classification.
[…]
@@ 3.3.1 Structured Debate and Judge
The CEJ process comprises four sequential prompt-based stages, designed to elicit, interrogate, and synthesize domain-relevant knowledge:
Initial Opinions. Each persona independently analyzes the input, providing an initial classification, justification, with a confidence score. This initial stage establishes diverse baseline perspectives.
Structured Debate. Personas are subsequently exposed to all initial opinions and begging a critical evaluation of their peers’ reasoning. During this deliberation, each persona must: (i) engage with at least one other perspective, agreeing or disagreeing while offering supporting rationale; (ii) revise their own stance if confronted with compelling counter-arguments; and (iii) reassess the input through the interpretive lens of other personas, re-evaluating both intent and target. This process results in a revised classification, an updated justification, and an adjusted confidence score for each persona.
Summarization. Then, a dedicated summarization agent condenses the full debate into a concise synthesis highlighting the main arguments, points of consensus, and unresolved disagreements.
Final Judgment. Finally, the judge model synthesizes all available evidence, i.e., the original input, both initial and revised persona opinions, and the debate summary, to produce a final adjudication. Prompted to act impartially, the judge delivers a final classification, comprehensive justification, and a confidence score.
   # Variant Model    Task A    Task B    Task C    Zero-shot Baselines    1 GPT-5.2    0.7306    0.4543    0.3381    2 Qwen-2.5-72B-Instruct    0.6190    0.4246    0.2576    3 LLaMA-3.3-70B-Instruct    0.6421    0.4297    0.2775    4 Cogito-70B    0.6106    0.4214    0.2683    5 Cogito-70B (reasoning)    0.6332    0.4325    0.2826    SemEval 2023 Task 10.A Leaderboard    6 DeBERTa-v3-large + twHIN-BERT-large Zhou (2023)    0.8746    –    –    7 RoBERTa-Large + ELECTRA    0.8740    0.7203    0.5487    8 DeBERTa Ensemble    0.8740    –    –    Data Augmentation and Ensemble Methods    9 SEFM Zhong et al. (2023)    0.8538    0.6619    0.4641    10 QCon Feely et al. (2023)    0.8400    0.6400    0.4700    11 HULAT Segura-Bedmar (2023)    0.8298    0.5877    0.4458    12 CSE Khan et al. (2025a)    0.8819    0.7243    0.5639    13 DDA Khan et al. (2025a)    0.8769    0.7277    0.6018    14 PaLM Ensemble    –    0.7326    –    Ours    15 $\mathcal{C}_{1}$ LLaMA-3.2-3B    0.7824    0.5941    0.3717    16 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific regulations    0.8986    0.6279    0.4189    17 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C    0.9195    0.7324    0.5842    18 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C    0.9214    0.7421    0.5904